Shamrock Industries uses process costing. All of the companys manufacturing activities take place in a single processing departm The following information was available for the month of June. Direct materials Direct labor Manufacturing overhead applied Total costs to account for in June 81,000 29,000 41,000 $151,000 The amounts of work in process at the beginning and end of the month were immaterial and assigned no dollar value. During June, 13,250 units were completed, of which 10,000 were sold on account at $25 per unit. a. Prepare a journal entry to summarize the total manufacturing costs applied to production in June. b. Prepare the journal entry to transfer completed units from work in process to the finished goods warehouse in June. c. Prepare the journal entries to record the sale of 10,000 units manufactured during the period and the related cost of goods sold. Complete this question by entering your answers in the tabs below Required A Required B Required C Prepare the journal entries to record the sale of 10,000 units manufactured during the period and the related cost of goods sold. (If no entry is required for a transaction/event, select No journal entry required in the first account field. Round your intermediate calculations to 2 decimal places.) K Prev 3 of 3 Next > e to search
Answer #1
Debit Credit
a Work in Process inventory 151000
        Direct materials 81000
        Direct labor 29000
        Manufacturing overhead 41000
b Finished goods inventory 151000
         Work in Process inventory 151000
c Accounts Receivable 250000 =10000*25
         Sales 250000
Cost of goods sold 114000 =10000*11.4
        Finished goods inventory 114000
Note: Cost per unit = 151000/13250=$11.40